OpenStudy (anonymous):

Carbon-14 has a half-life of 5,730 years. Suppose a fossil from a mammal contains one-fourth of its original amount of carbon-14. How old is the sample? 22,920 years 17,190 years 2,865 years 11,460 years

OpenStudy (anonymous):

i think its 22920 years ?

OpenStudy (anonymous):

no/n1=1/2t/1/2lif

OpenStudy (anonymous):

what does that mean ..

OpenStudy (aaronq):

5730 = ln2/k k=1.2x10^-4 A=Aoe^(-kt) 25=100e^(-1.2x10^-4t) ln(.25)=-1.2x10^-4t t=-1.3862/-1.2x10^-4= 11552.45 years old It's a bit off cuz i rounded

OpenStudy (anonymous):

so it would be d